A boy,pulls a 1.0 kg toy horizontally in a straight line at a constant speed of 20 m/min by exerting a 3.0 N force at an angle of 30 degrees with the horizontal.What is the average power developed by the boy?
1
Expert's answer
2010-11-04T14:41:55-0400
The average power will be equal to the total work (F*S, there F is the force, S - is the path) per some time (t):

W = (F*S)/t = F * V (v - speed);

As the force exerts at the angle to the horizont, we have to take its projectile:

F = F' cos (30) = 3 * 0.867 = 2.6




20 m/min = 20/60 = 0.33 m/sec.

W = 2.6 * 0.33 = 0.86 [W].